A population of values has a normal distribution with μ = 149.8 and σ = 25.6 . You intend to draw a random sample of size n = 103 . Find the probability that a single randomly selected value is between 148.3 and 157.6. P(148.3 < X < 157.6) = 0.094 Incorrect Find the probability that a sample of size n = 103 is randomly selected with a mean between 148.3 and 157.6. P(148.3 < M < 157.6) = Incorrect Enter your answers as numbers accurate to 4 decimal places. Answers obtained using exact z-scores or z-scores rounded to 3 decimal places are accepted.
Answer #1

Answer

we have μ = 149.8, σ 25.6

(a) We have to find the probability that a single randomly selected value is between 148.3 and 157.6

so, x1 = 148.3 and x2 = 157.6

we have to find the value of P(x_1<x<x_2)

using the formula P(x_1<x<x_2) = P((ar{x_1}-mu)/sigma<z<(ar{x_2}-mu)/sigma)

setting the given values, we get

P(1483 < x < 157.6) = P(048.3-149.8)/25.6 < z < (157.6-149.8)/256)

this gives us

P(148.3 < x < 157.6) = P(-0.0586 < z < 0.3047)

we can write it as

P(148.3 < x < 157.6) = P(z < 0.3047)-P(z <-0.0586)

using the z distribution table, we get

P( 148.3 < x < 157.6) 0.6179-0.4761-0.1418

So, the required probability is 0.1418

(2)

We have to find the probability that a single randomly selected value is between 148.3 and 157.6

so, x1 = 148.3 and x2 = 157.6 and sample size is n = 103

we have to find the value of P(x_1<x<x_2)

using the formula P(x_1<x<x_2) = P((ar{x_1}-mu)/(sigma/sqrt{n})<z<(ar{x_2}-mu)/(sigma/sqrt{n}))

setting the given values, we get

P(148.3 <r< 157.6) P((148.3-149.8)/(25.6/103)z (157.6- 149.8)/(25.6/V103))

this gives us

P(1483 < x < 157.6)-Pl (-1.5/2.52) <之< (7.8/2.52))

this gives us

Pl 148.3 < x < 157.6) = P(-0.595 < z < 3.092)

we can write it as

P(1483 < x < 157.6) = P(z < 3.092)-P(z <-0.595)

using the z distribution table, we get

P(148.3 < x < 157.6) = 0.999-0.2743 = 0.7247

So, the required probability is 0.7247